# Calibrating Long-form Generations from Large Language Models

### v1

@@ 5.1 Correlation
We assess the dataset-level calibration by measuring the correlation between expected values of confidence and correctness distributions. For each answer $A_{i}$ within the dataset, we can calculate the expected confidence score $E(C_{i})$ and the expected correctness score $E(G_{i})$. In the whole dataset, we can get a list of expected confidence scores $\mathbf{E_{C}}$ and a list of expected correctness scores $\mathbf{E_{G}}$. Then we can measure the correlation between them.
$\rho(\mathcal{D})=Corr(\mathbf{E_{C}},\mathbf{E_{G}})$ (6)
@@
where $Corr(\cdot)$ represents the correlation function (Pearson correlation in our experiments). This correlation provides a clear indicator of how well our model’s confidence aligns with its actual correctness across the entire dataset.
5.2 Wasserstein Similarity
Wasserstein distance represents how much “work” it takes to transform one distribution into another. Wasserstein similarity $S_{W}$ adapts the concept of Wasserstein distance $D_{W}$ in a way that it represents similarity rather than distance between two probability distributions. Formally, the instance-level Wasserstein similarity for an answer $A_{i}$ is:
$S_{W}(A_{i})=1-\frac{D_{W}(G_{i},C_{i})}{d_{max}}$ (7)
where $d_{max}$ is the maximum distance in the metric space. And the dataset-level calibration score can be obtained by averaging the instance-level calibration score.
$S_{W}(\mathcal{D})=\frac{1}{|\mathcal{D}|}\sum_{i}{S_{W}(A_{i})}$ (8)
5.3 Selective F1
In selective answering (e.g., Kamath et al., 2020), models only respond when confident about their accuracy, enhancing their reliability. To assess the effectiveness of LMs in selective answering, we propose the selective F1 metric ($F1_{\tau_{s}}$). This metric quantifies the model’s aptitude in filtering out answers that meet or exceed a predefined correctness threshold, denoted as $\tau_{s}$. Alongside this, we define two critical components: selective precision ($P_{\tau_{s}}$) and selective recall ($R_{\tau_{s}}$). Selective precision calculates the proportion of selected answers that surpass the correctness threshold $\tau_{s}$ relative to the total number of selected answers. In contrast, selective recall compares the number of selected answers meeting this criterion against the total number of correct answers in the dataset that exceed the threshold $\tau_{s}$.
@@ 5.3 Selective F1
Our approach utilizes a dual-threshold system, consisting of a confidence threshold ($\tau_{c}$) and the correctness threshold ($\tau_{s}$), allowing the model to answer questions only if its confidence in the answer’s expected correctness score exceeding $\tau_{s}$ surpasses $\tau_{c}$. To more formally define the selective F1, let $\mathcal{A}=\{A_{1},A_{2},...,A_{n}\}$ denote the total set of model’s answers in the dataset and $\mathcal{A}^{*}=\{A_{i}\in\mathcal{A}\mid\text{if }\sum_{s\geq\tau_{s}}C_{i}(s)\geq\tau_{c}\}$ denote the set of selected answers. Let the indicator function $I_{\tau_{s}}(A_{i})$ indicate if the expected correctness score $E[G_{i}]$ of $A_{i}$ exceeds $\tau_{s}$
$I_{\tau_{s}}(A_{i})=\begin{cases}1&\text{if }E[G_{i}]\geq\tau_{s}\\ 0&\text{Otherwise}\end{cases}$ (9)
The selective precision on the dataset $\mathcal{D}$ is:
@@
The selective recall is:
@@
The selective F1 takes both recall and precision into account:
@@
ASQA ELI5
Model Method Corr Wass $F1_{0.8}$ Score Corr Wass $F1_{0.8}$ Score
Llama2-13b Self Eval 18.1 71.5 51.0 51.3 21.8 75.0 39.6 53.8
Self Cons 48.0 75.8 3.6 28.8 61.0 39.8
Llama2-70b Self Eval 16.4 64.9 60.3 59.4 18.8 78.9 54.9 61.7
Self Cons 44.3 77.0 29.7 15.5 71.4 51.4
Vicuna-13b Self Eval 30.2 70.8 30.0 50.8 31.2 76.6 36.7 53.0
Self Cons 58.2 72.0 37.6 35.3 63.8 36.9
GPT-3.5-turbo Self Eval 3.6 74.3 83.7 72.6 26.2 76.9 63.5 63.0
Self Cons 30.5 77.7 58.4 18.5 68.6 55.2
QAMPARI CNNDM
Model Method Corr Wass $F1_{0.4}$ Score Corr Wass $F1_{0.8}$ Score
Llama2-13b Self Eval 25.7 66.6 29.0 13.3 19.5 96.5 92.0 77.0
Self Cons 46.0 71.6 42.3 9.6 73.2 24.5
Llama2-70b Self Eval 18.6 41.9 30.1 14.6 6.4 96.9 93.6 77.6
Self Cons 17.1 61.0 39.1 22.7 72.4 20.2
Vicuna-13b Self Eval 16.2 62.6 27.0 11.4 3.9 80.9 64.2 78.2
Self Cons 49.5 78.7 39.7 18.9 78.5 41.7
GPT-3.5-turbo Self Eval 28.7 57.6 53.1 24.0 17.0 96.2 94.7 78.2
Self Cons 44.4 65.8 56.0 18.9 77.8 32.1
Table 1: Calibration Performance Across Four Tasks (in %): “Corr” for Correlation, “Wass” for Wasserstein Similarity, “$F1_{\tau_{s}}$” for Selective F1 Score at threshold $\tau_{s}$, and “Score” representing expected correctness. “Self Eval” for self-evaluation and “Self Cons” for self-consistency. Bold indicates the best score. Key insights include: 1) Higher model capability does not ensure better calibration; 2) Models may exhibit calibration on certain metrics but not others; 3) On factoid datasets (ASQA, QAMPARI), self-consistency outperforms self-evaluation.
@@ 6.1 Setup
Models and Data We measure different sized LLMs’ calibration, including Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), vicuna-13b (Zheng et al., 2023), GPT-3.5-turbo, acorss three long-form QA tasks: ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022), and one summarization task: CNNDM (Nallapati et al., 2016). Details of each dataset can be found in Appendix A.
Correctness Evaluation We apply GPT-4 to evaluate the model-generated answers to approximate the ground-truth correctness distributions for ASQA, ELI5, and CNNDM. In QAMPARI, given that the answer is a list of entities, we evaluate using the F1-5 metric (Gao et al., 2023), calculating the F1 score by the exact match with the gold answer list and defining 100% recall for predictions with at least 5 correct answers. In preliminary experiments, we found that samples from the GPT-4 evaluator had very low variance in terms of the final score assigned to the answers, hence in the following results we only use a single sample, resulting in a point mass distribution for correctness.
Confidence Elicitation We contrast self-evaluation and self-consistency methods for confidence assessment across various tasks, employing task-specific strategies to compare similarities. For ASQA, where answers hinge on factual claims, we utilize a self-consistency approach focused on claims. In the case of QAMPARI, which involves listing entities, we adopt self-consistency-NER. For the ELI5 dataset, a more general self-consistency-naive strategy is applied due to the broad nature of the questions. Lastly, for CNNDM, where summaries are composed of key sentences, we implement a self-consistency-split for detailed analysis. These tailored approaches ensure the most effective evaluation of each dataset’s unique characteristics.
@@ 6.2 Main Results
Larger models are not necessarily better calibrated. Table 1 reveals a surprising trend: Vicuna-13b, though performing poorly on ASQA, ELI5 and QAMPARI, has the highest correlation with self-consistency confidence on the same datasets. This might be due to the fact that Vicuna-13b is the only model instruction fine-tuned in a supervised fashion, in contrast to the other models using reinforcement learning human feedback. RL finetuning has been shown to lead to miscalibration in LMs by biasing generations towards highly rewarded behaviors (Kadavath et al., 2022). Scaling the temperature could enhance the calibration of LLMs fine-tuned using RL (see § 6.3). Additionally, Llama-2-13b demonstrates a higher correlation than its larger counterpart, Llama-2-70b, in $3$ out of $4$ datasets with both confidence elicitation methods. However, GPT-3.5-turbo, the strongest model, consistently scores the highest in selective F1 across all datasets. Selective F1 metrics blend performance and calibration and tend to yield higher scores for stronger models.
A model can be calibrated on some dimensions but not on others. Calibration metrics offer complementary perspectives. A model may exhibit low correlation but high Wasserstein similarity, or vice versa. For instance, GPT-3.5-turbo on the ASQA dataset shows low correlation ($3.6$) but high Wasserstein similarity ($74.3$) with self-evaluation confidence. In contrast, Vicuna-13b demonstrates the opposite trend on ASQA with self-consistency confidence (correlation: $58.2$, Wasserstein similarity: $72.0$). This discrepancy arises because correlation and Wasserstein similarity capture different calibration aspects: correlation reflects the model’s ability to appropriately rank good and bad answers, while Wasserstein similarity evaluates how closely a model’s confidence mirrors actual correctness.
Self-Consistency is better than Self-evaluation on factoid datasets. In factoid datasets (ASQA and QAMPARI), self-consistency confidence elicitation outperforms self-evaluation for all models. On QAMPARI, all models show better calibration with self-consistency across all four metrics, while on ASQA, the trend holds for all but selective F1. While self-consistency scores can serve as indicators of answer quality, they are not guaranteed to be the same as ground-truth correctness due to the inability to fully grasp the ground-truth criteria. However, the effectiveness of self-consistency diminishes in more open-ended tasks like ELI5 or the CNNDM summarization dataset. As tasks become more subjective, the power of self-consistency in model calibration begins to wane.
@@ 6.3 Improving Calibration
We delve into four strategies aimed at enhancing the calibration score of LLMs: fine-tuning the model to do self-evaluation, retrieving source documents, scaling the temperature, and hybrid confidence elicitation strategies.
Fine-tuning Our study examines enhancing model calibration through fine-tuning, which is conducted on the ASQA dataset. We generate a self-evaluation training dataset using GPT-4 to assess answers from various models to ASQA’s training set questions. This dataset comprises inputs of questions and corresponding model answers, with outputs including a score and an explanation for each answer. The Llama2-13b model is then fine-tuned using LoRA (Hu et al., 2021) on this dataset, aiming to improve its scoring accuracy and explanatory capabilities. We also consider the impact of self-evaluation finetuning on the model’s ability to generate responses, and further experiment with finetuning the model on a combined objective for self-evaluation as well as generating answers from the ASQA training set. For comparison, we also fine-tune the model using generation data only. See Appendix D.1 for details of the training methodology. As Table 2 reveals, solely training on self-evaluation (‘Evaluation’) did not yield consistent improvements in calibration, possibly due to the complexity of this task, as well as the limitation of LORA fine-tuning. Nonetheless, fine-tuning the model improves the self-consistency method, especially when the generation data is included during training (‘Eval + Gen’ and ‘Generation’). The model becomes more confident in terms of self-consistency after fine-tuning.
[…]
Self-Evaluation
None 18.1 $\uparrow$ 71.5 $\uparrow$ 51.0 $\uparrow$ 51.3 $\uparrow$
Evaluation 13.6 $\downarrow$ 67.0 $\downarrow$ 62.0 $\uparrow$ 49.2 $\downarrow$
Generation 20.0 $\uparrow$ 72.4 $\uparrow$ 52.8 $\uparrow$ 52.1 $\uparrow$
Eval + Gen 23.9 $\uparrow$ 74.2 $\uparrow$ 46.6 $\downarrow$ 50.1 $\downarrow$
Self-Consistency
None 48.0 $\uparrow$ 75.8 $\uparrow$ 3.6 $\uparrow$ 51.3 $\uparrow$
Evaluation 46.9 $\downarrow$ 76.7 $\uparrow$ 10.7 $\uparrow$ 49.2 $\downarrow$
Generation 58.9 $\uparrow$ 76.4 $\uparrow$ 32.4 $\uparrow$ 52.1 $\uparrow$
Eval + Gen 54.5 $\uparrow$ 78.4 $\uparrow$ 18.9 $\uparrow$ 50.1 $\downarrow$
@@
Source Documents We examine how additional contextual information influences LLM calibration. Using the Llama-2-13b-chat and GPT-3.5-turbo on ASQA dataset, with a focus on self-consistency as our confidence eliciation method, we test the model’s performance when supplemented with different types of source documents. We experiment with two types of documents: random documents related to the question and ‘oracle’ documents directly relevant to the answers, as included in the dataset release Stelmakh et al. (2022). Our findings, illustrated in Table 3, reveal that oracle documents can enhance model performance and calibration across two out of three metrics for both models, while random documents may be less effective. These results underscore the importance of relevant contextual information in model calibration.
[…]
R 39.0 $\downarrow$ 75.6 $\downarrow$ 30.6 $\uparrow$ 59.6 $\uparrow$
O 49.5 $\uparrow$ 75.6 $\downarrow$ 20.1 $\uparrow$ 65.8 $\uparrow$
[…]
R 36.6 $\uparrow$ 71.8 $\downarrow$ 68.2 $\uparrow$ 64.8 $\downarrow$
O 27.0 $\downarrow$ 79.5 $\uparrow$ 72.8 $\uparrow$ 75.8 $\uparrow$
Table 3: How source documents affect the calibration score. In the document column, “N” means no documents in the input prompt, “R” means randomly selected documents relevant to the topic, and “O” means the oracle documents relevant to the answer. $\uparrow$ denotes that the calibration score goes up when adding documents, while $\downarrow$ means going down.
Temperature We adjust the generation temperature for Llama2-13b from 0.2 to 1 to examine its impact on calibration. The result in Figure 3 reveals a consistent upward trend in all calibration metrics. Notably, the model’s performance initially improves and then deteriorates. This observation implies that modulating the generation temperature can enhance the calibration of the reinforcement-learned model, which is consistent with previous findings observed in short-form answer generation (Kadavath et al., 2022).
@@
Hybrid Confident Elicitation Prior work (Xiong et al., 2023, Tian et al., 2023b) found that combining self-evaluation and self-consistency can yield a more accurate confidence score. In our study, we explore whether this finding applies to long-form generation. Specifically, for an answer $A_{i}$, we merge confidence distributions obtained through both self-evaluation ($C^{eval}_{i}$) and self-consistency ($C^{consist}_{i}$). The resultant hybrid confidence distribution, $C^{hybrid}_{i}=\alpha C^{eval}_{i}+(1-\alpha)C^{consist}_{i}$, is modulated by the parameter $\alpha$, allowing us to adjust the influence of each distribution. Our experiments are conducted on the ASQA dataset and utilize the correlation as the calibration metric. As shown in Figure 4, we observe that the correlation between confidence and correctness initially increases but then declines as $\alpha$ is varied from 0 to 1. However, this trend is not observed in other metrics (See Figure 7), indicating that while hybrid calibration elicitation may enhance calibration in terms of correlation, it may not have the same impact on other dimensions.
[…]
Figure 4: Correlation vs. Weight Variation for hybrid confidence elicitation. $\alpha$ is the weight for self-evaluation confidence. $\alpha=0$ for solely self-consistency and $\alpha=1$ for solely self-evaluation.

### v2

@@ 3.4 Calibration Metrics
Correlation While ECE-M focuses on measuring calibration at each correctness level independently, it doesn’t account for the distance between these levels. To address this, we also measure the correlation between the expected values of the confidence and correctness distributions for a more comprehensive assessment. For each answer $A_{i}$ within the dataset, we can calculate the expected correctness score of the model $E|C_{i}|=\sum_{s}P_{C_{i}}(s)\times s$ and the expected correctness score of the target $E|T_{i}|=\sum_{s}P_{T_{i}}(s)\times s$. In the whole dataset, we can get a list of expected correctness scores $\mathbf{E_{C}}$ from model confidence and a list of target expected correctness scores $\mathbf{E_{T}}$. Then we can measure the correlation between them.
$\rho(\mathcal{D})=Corr(\mathbf{E_{C}},\mathbf{E_{T}})$ (7)
@@
where $Corr(\cdot)$ represents the correlation function. This correlation provides a clear indicator of how well our model’s confidence aligns with its actual correctness across the entire dataset.
Selective F1 In selective answering (e.g., Kamath et al., 2020), models only respond when confident about their accuracy to improve reliability. Traditional metrics for selective answering include accuracy@coverage and coverage@accuracy (Tian et al., 2023b), which measures a model’s precision and recall in selecting completely correct answers. Similarly, in long-form calibration, it is crucial to assess the model’s selection of answers that are at least "scorrect" using both precision and recall.
@@
Therefore, we propose the selective F1 metric ($F1_{\tau_{s}}$) to quantify the model’s aptitude in filtering out answers that meet or exceed a predefined correctness threshold $\tau_{s}$. Our approach utilizes a dual-threshold system, consisting of a confidence threshold ($\tau_{c}$) and the correctness threshold ($\tau_{s}$), allowing the model to answer questions only if its confidence in the answer’s expected correctness score exceeding $\tau_{s}$ surpasses $\tau_{c}$. Formally, let $\mathcal{A}=\{A_{1},A_{2},...,A_{n}\}$ denote the total set of model’s answers in the dataset and $\mathcal{A}^{*}=\{A_{i}\in\mathcal{A}\mid\sum_{s\geq\tau_{s}}P_{C_{i}}(s)\geq\tau_{c}\}$ denote the set of selected answers. Let the indicator function $I_{\tau_{s}}(A_{i})$ indicate if the expected correctness score $E[T_{i}]$ of $A_{i}$ exceeds $\tau_{s}$:
$I_{\tau_{s}}(A_{i})=\begin{cases}1&\text{if }E[T_{i}]\geq\tau_{s}\\ 0&\text{Otherwise}\end{cases}$ (8)
The selective precision $P_{\tau_{s}}$ on the dataset $\mathcal{D}$ is the proportion of selected answers that surpass the correctness threshold $\tau_{s}$ relative to the total number of selected answers:
@@
The selective recall $R_{\tau_{s}}$ compares the number of selected answers meeting this criterion against the total number of correct answers in the dataset that exceed the threshold $\tau_{s}$:
@@
The selective F1 combines recall and precision:
@@
In our experiments, we select $\tau_{s}$ as the nearest correctness level greater than the best LM’s average correctness score. For $\tau_{c}$, we choose the value that yields the highest selective F1 score on the development split.
ASQA QAMPARI ELI5 CNNDM
Method ECE-M Corr $F1_{0.8}$ ECE-M Corr $F1_{0.4}$ ECE-M Corr $F1_{0.8}$ ECE-M Corr $F1_{0.8}$
SL 28.2$\uparrow$ 0.7$\uparrow$ 0.0 $\uparrow$ 27.4$\uparrow$ 10.9$\uparrow$ 8.0$\uparrow$ 29.6$\uparrow$ -11.9$\uparrow$ 0.0$\uparrow$ 77.2$\uparrow$ -7.5$\uparrow$ 0.0$\uparrow$
BSE 32.8$\uparrow$ 14.2$\uparrow$ 57.6 $\uparrow$ 25.2$\uparrow$ 16.8$\uparrow$ 33.2$\uparrow$ 30.3$\uparrow$ 11.7$\uparrow$ 46.7$\uparrow$ 78.5$\uparrow$ 11.2$\uparrow$ 90.6$\uparrow$
CSE 29.0$\downarrow$ 16.3$\uparrow$ 58.5 $\uparrow$ 42.8$\uparrow$ 21.9$\uparrow$ 33.4$\uparrow$ 31.2$\uparrow$ 26.9$\uparrow$ 48.2$\uparrow$ 15.2$\downarrow$ 19.2$\uparrow$ 92.0$\uparrow$
ASC 35.9$\uparrow$ 27.1$\uparrow$ 5.2 $\uparrow$ 46.0$\uparrow$ 38.6$\uparrow$ 38.5$\uparrow$ 38.4$\uparrow$ 16.7$\uparrow$ 7.9$\uparrow$ 63.2$\uparrow$ 8.8 $\uparrow$ 44.2$\uparrow$
PSCF1 28.8$\uparrow$ 27.1$\uparrow$ 33.5 $\uparrow$ 38.4$\uparrow$ 38.6$\uparrow$ 42.7$\uparrow$ 27.1$\uparrow$ 16.7$\uparrow$ 20.7$\uparrow$ 57.1$\uparrow$ 8.8$\uparrow$ 79.5$\uparrow$
PSC 18.3$\downarrow$ 46.8$\uparrow$ 61.6 $\uparrow$ 26.2$\downarrow$ 39.1$\uparrow$ 44.0$\uparrow$ 24.9$\downarrow$ 24.9$\uparrow$ 46.2$\uparrow$ 64.5$\uparrow$ 15.5$\uparrow$ 90.0$\uparrow$
Table 1: Calibration Performance Comparison Among Different Confidence Elicitation Methods Across Four Tasks (in %): “ECE-M” for expected correctness error with multi-class, “Corr” for Correlation, “$F1_{\tau_{s}}$” for Selective F1 Score at threshold $\tau_{s}$. Results represent averages from five models. Methods with * served as baselines. For “Corr”, “$F1_{\tau_{s}}$”, and “Score”, $\uparrow$ means better than corresponding baseline while $\downarrow$ is worse. For “ECE-M”, $\downarrow$ is better while $\uparrow$ is worse. The best score among all confidence elicitation methods is bolded. Key insights: 1) Self-Consistency (PSC) outperforms Self-Evaluation on factoid datasets; 2) Our methods PSC and CSE surpass baselines; 3) Different metrics offer complementary insights
@@ 4.1 Setup
Models and Data We measure different sized LLMs’ calibration, including Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), Vicuna-13b (Zheng et al., 2023), Llama-3-8b-Instruct, GPT-3.5-turbo, across three long-form QA tasks: ASQA (Stelmakh et al., 2022), ELI5 (Fan et al., 2019), QAMPARI (Amouyal et al., 2022), and one summarization task: CNNDM (Nallapati et al., 2016). Details of datasets can be found in Appendix A.
Correctness Evaluation We apply GPT-4 to evaluate target correctness distributions for ASQA, ELI5, and CNNDM. In QAMPARI where the answer is a list of entities, we evaluate using the F1-5 metric, calculating the F1 score by the exact match with the gold answer and defining 100% recall for predictions with at least 5 correct answers.
4.2 Confidence Elicitation Methods
In addition to our methods CSE and PSC (see Appendix D.3 for similarity measurement choosing), we established baselines for self-evaluation, self-consistency, and logits-based approaches. This is because prior studies lack directly applicable baselines, primarily due to the non-interpretable nature of single confidence scores.
Sentence Likelihood (SL): Based on prior studies using logits to gauge model confidence, we adopt sentence likelihood as a baseline measure, which typically results in a confidence distribution focused at the lowest score in long-form answers.
Binary Self-Evaluation (BSE): following previous work Kadavath et al. (2022) that asks model to self-evaluate if its answer is true several times, using the frequency of true as model’s confidence score towards the answer being true. Then we adapt such a single score as a distribution focus solely on the values 0 and 1.
Average Self-Consistency (ASC): Following prior work Xiong et al. (2023) using the average consistency between these candidate responses and the original answer then serves as a single measure of confidence score, we adopt the simple f1-token score to measure the consistency to adapt it to long-form generations. Then we treat the single score as a point mass distribution.
Pairwise Self-Consistency F1 (PSC-F1): Still using F1 to measure the consistency like ASC, but we directly treat the pairwise consistency scores as a distribution without aggregating, thereby keeping model intrinsic understanding about different correctness levels.
@@ 4.3 Main Results
In Table 1, we evaluate the calibration performances of various confidence elicitation methods by averaging the scores across all models. Table 1 shows that our methods, CSE and PSC, generally outperform their respective baseline categories and also surpass the logits-based method SL. Key findings from the results include:
Self-Consistency Outperforms Self-Evaluation on Factoid Datasets Self-consistency methods typically outperform self-evaluation on factoid datasets like ASQA and QAMPARI. However, their effectiveness diminishes in more subjective tasks such as ELI5 or CNNDM. We hypothesize that this is because self-consistency is more readily quantifiable in factoid datasets, where the agreement between answers can be assessed based on factual consistency, thus providing clearer criteria. Conversely, in open-ended datasets, the consistency between answers is more ambiguous, making it more difficult to measure.
Nuanced Self-Evaluation Enhances Calibration CSE generally outperforms BSE by providing detailed confidence estimates at each correctness level. However, the overall improvement remains constrained by the intrinsic limitations of LMs in self-evaluating their correctness, which sometimes hampers accurate estimations.
Pair-wise Similarities Distribution and Task-tailored Similarity Measurement Help Calibration Both ASC and PSC-F1 measure similarity with token-level F1. However, PSC-F1 treats these scores as a distribution rather than averaging them, leading to better ECE-M and selective F1. PSC further enhances calibration by adopting a task-specific, detailed measurement of similarity, outperforming FSC in all four tasks.
ASQA QAMPARI
Model ECE-M Corr $F1_{0.8}$ Score ECE-M Corr $F1_{0.4}$ Score
Llama2-13b 15.9 48.0 47.5 51.3 30.5 46.0 42.3 13.3
Llama2-70b 14.7 44.3 61.9 59.4 29.3 17.1 37.0 14.6
Vicuna-13b 20.1 58.2 56.8 50.8 14.4 49.5 42.6 11.4
Llama-3-8b 14.3 53.2 65.1 54.9 33.0 38.3 42.0 14.1
GPT-3.5-turbo 26.7 30.5 76.7 72.6 23.7 44.4 56.2 24.0
ELI5 CNNDM
Model ECE-M Corr $F1_{0.8}$ Score ECE-M Corr $F1_{0.8}$ Score
Llama2-13b 36.0 21.8 40.1 53.8 12.6 19.5 92.0 77.0
Llama2-70b 32.8 18.8 54.4 61.7 13.8 6.4 93.6 77.6
Vicuna-13b 19.4 31.3 34.9 53.0 32.7 3.9 64.2 78.2
Llama-3-8b 34.9 36.7 48.4 57.1 9.3 49.0 86.9 77.8
GPT-3.5-turbo 32.7 26.2 63.4 63.0 7.8 17.0 94.7 78.2
Table 2: Comparison of Calibration Performance Across Models for Four Tasks (in %): We identify the optimal confidence elicitation method for each task and compare the performance of various models using this method. “Score” means the model’s average correctness score on that task. A key observation is that more powerful LMs do not necessarily exhibit better calibration, although they tend to perform better in selective answering.
Calibration metrics complement each other A simplistic approach like SL, which allocates all the probability mass to the point of score 0, can misleadingly show decent ECE-M (28.2%) in specific cases like ASQA. However, its negative correlation (0.7%) and zero F10.8 underscore an ineffective confidence distribution. Similarly, BSE in CNNDM may achieve a high F10.8 (90.6%) by overestimating answer correctness, but this does not truly reflect response quality (correlation: 11.2%) or provide well-calibrated probabilities across correctness levels, resulting in a bad ECE-M (ECE-M: 78.5%). Hence, a comprehensive evaluation using multiple metrics is essential for a balanced assessment of model calibration.
@@
Larger models are not necessarily better calibrated. In Table 2, we focus on the calibration performance of individual models when paired with the best-performing confidence elicitation method for each task. Table 2 shows that despite poor performance on ASQA and QAMPARI, Vicuna-13b has the highest correlation across these datasets. It might be because reinforcement learning for other models causes miscalibration by encouraging overfitting to rewarded behaviors (Kadavath et al., 2022). Scaling the temperature could enhance the calibration of LLMs fine-tuned using RL (see § 4.4). Additionally, Llama-2-13b demonstrates a higher correlation than its larger counterpart, Llama-2-70b. However, GPT-3.5-turbo, the strongest model, consistently scores the highest in selective F1 across all datasets. This performance can be attributed to the model’s ability to generate a larger volume of high-quality answers, increasing the probability of selecting superior responses even if it is not particularly well-calibrated. Consequently, the Selective F1 metric blends performance and calibration, and tends to favor more capable models due to their higher output of quality answers.
@@ 4.4 Improving Calibration
We delve into different strategies to enhance calibration: fine-tuning, scaling the temperature, adding source documents (Appendix E.2), and hybrid confidence elicitation (Appendix E.3).
Fine-tuning Our study explores three fine-tuning strategies to improve model calibration on the ASQA dataset: fine-tuning the model for self-evaluation (using questions and model answers to produce scores and explanations), fine-tuning the model for generation (generating answers from questions), and a hybrid of both. GPT-4 synthesizes self-evaluation data by assessing different models’ answers to questions from the ASQA training set, while the generation data originates from the ASQA training set itself. We apply LoRA (Hu et al., 2021) fine-tuning to the Llama2-13b model. See Appendix E.1 for experiment details. As Table 3 reveals, solely training on self-evaluation (‘Evaluation’) did not yield consistent improvements in calibration, possibly due to the complexity of this task and the limitation of LORA. Nonetheless, fine-tuning the model improves the self-consistency method, especially when the generation data is included during training (‘Eval + Gen’ and ‘Generation’). The model becomes more confident in terms of self-consistency after fine-tuning.
Data Corr ECE-M F10.8 Score
Self-Evaluation (CSE)
None 18.1 $\uparrow$ 30.3 $\uparrow$ 50.8 $\uparrow$ 51.3 $\uparrow$
Evaluation 13.6 $\downarrow$ 32.4 $\uparrow$ 52.3 $\uparrow$ 49.2 $\downarrow$
Generation 20.0 $\uparrow$ 26.2 $\downarrow$ 53.4 $\uparrow$ 52.1 $\uparrow$
Eval + Gen 23.9 $\uparrow$ 20.2 $\downarrow$ 46.6 $\downarrow$ 50.1 $\downarrow$
Self-Consistency (PSC)
None 48.0 $\uparrow$ 15.9 $\uparrow$ 47.5 $\uparrow$ 51.3 $\uparrow$
Evaluation 46.9 $\downarrow$ 13.6 $\downarrow$ 56.1 $\uparrow$ 49.2 $\downarrow$
Generation 58.9 $\uparrow$ 14.5 $\downarrow$ 59.5 $\uparrow$ 52.1 $\uparrow$
Eval + Gen 54.5 $\uparrow$ 12.2 $\downarrow$ 50.7 $\uparrow$ 50.1 $\downarrow$
@@
Temperature We adjust the generation temperature for Llama2-13b from 0.2 to 1 to examine its impact on calibration. The result in Figure 3 reveals consistent improvements in all calibration metrics. Notably, the model’s performance initially improves and then deteriorates. This observation implies that modulating the generation temperature can enhance the calibration of the model.